Melodyis three times as old as jaypee.Ten years ago she was eight times as old as jaypee. How old are they?
Answer by Cromlix(4381)

Hi there,
Melodyis three times as old as jaypee.
M = 3J
Ten years ago she was eight times as old as jaypee.
M - 10 = 8(J - 10)
Replace M with 3J from top equation.
3J - 10 = 8(J - 10)
3J - 10 = 8J - 80
Collect like terms.
3J - 8J = -80 + 10
-5J = -70
5J = 70
J = 14
Melodyis is 42 years old
Jaypee is 14 years old.
Hope this helps :-)